How can I solve the word problem below, could you also go over how to understand what the words are explaining to me? Like, how to get the equations I need using the information in the question to answer it.
"The sum of two numbers is 56. The larger exceeds twice the smaller by 2. Find the numbers."

Answer by ankor@dixie-net.com(22740)

let the two number be a & b, write an equation for each statement:
:
""The sum of two numbers is 56."
a + b = 56
" The larger exceeds twice the smaller by 2."
a = b + 2
use the above equation to replace a in the 1st equation
(b+2) + b = 56
2b + 2 = 56
2b = 56 - 2
2b = 54
b = 54/2
b = 27
then
a = 27 + 2
a = 29
:
Check these values for a & b in both statements, see if they are true

Find the numbers."